please draw the ray diagrams for me as well

media%2Fb24%2Fb248ebe8-0565-4c02-8fbf-5d

please draw the ray diagrams for me as well
Answer #1

the thin lens equation gives image position for first lens

Q1=  P1F1 / P1-F1

Q1= 30*15 / 30-15

Q1= 30 cm

magnification by this lens M1= - Q1/ P1

M1= -30/30 = -1.00 cm

the real image formed by the first lens.

for second lens,

P2= 20- Q1= 20-30= -10 cm

Q2= P2*F2 / P2-F2

Q2= -10*15 / -10-15

Q2= 6cm

magnification by the second lens

M2= -Q2 /P2

M2= -6/-10

M2= 0.6 cm

virtual image is located by second lens.

the overall magnification M= M1*M2

M= (-1.0)*(0.6)

M= -(0.6)